The population standard deviation for the height of college basketball players is 3.4 inches. If we want to estimate 99% confidence interval for the population mean height of these players with a 0.43 margin of error, how many randomly selected players must be surveyed?
Answer #1

Solution :

Given that,

standard deviation = =3.4

Margin of error = E = 0.43

At 99% confidence level the z is,

= 1 - 99%

= 1 - 0.99 = 0.01

/2 = 0.005

Z/2 = 2.58 ( Using z table ( see the 0.005 value in standard normal (z) table corresponding z value is 2.58 )  

sample size = n = [Z/2* / E] 2

n = ( 2.58*3.4 / 0.43 )2

n =416.16

Sample size = n =416